OpenStudy (anonymous):

I WILL MEDAL!! and fan help please ANYONE!!!! How many solutions does the equation 5m − 5m − 12 = 14 − 2 have?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

explanation please!

OpenStudy (anonymous):

coming

OpenStudy (anonymous):

thanks dude!

OpenStudy (anonymous):

5m-5m-12=14-2 \mathrm{Add\:similar\:elements:}\:5m-5m=0 =0-12 \:0-12=-12 =-12 =-12 -12+12=12+12 0=24 false for all m

OpenStudy (anonymous):

hurry lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

more info

OpenStudy (anonymous):

so no solution

OpenStudy (anonymous):

yah none

OpenStudy (anonymous):

thanks bye

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

welcome